OpenStudy (anonymous):

Solve the logarthmic equation. 2 log3(x + 4) - log39 = 2

OpenStudy (gorv):

\[2*\log_3(x+4)-\log_3 9=2\]

OpenStudy (gorv):

\[c*\log_a b=\log_ab^c\]

OpenStudy (gorv):

or 9=3^2

OpenStudy (gorv):

\[\log_3 9=\log_3 3^2=2*\log_3 3\]

OpenStudy (gorv):

log3 3 =1

OpenStudy (gorv):

\[2*\log_3 (x+3)-2*\log_3 3 =2\]

OpenStudy (gorv):

divide each term by 2

OpenStudy (gorv):

what u got ??